OpenStudy (anonymous):

ok if my points are (y108,000, x2.0) and (y96000, x4.0) I need to find the slope of the line and put it into an equation.

OpenStudy (anonymous):

Slope : \[\frac{f(x) - f(x')}{x-x'}\] Just fill the correct value given in this generic formula.

OpenStudy (anonymous):

Once the slope in found, your (assumed linear) function will be : \[f(x) = \text{slope} \times x + \text{offset at (0,y)}\]

OpenStudy (anonymous):

To find the offset, just use solve this simple equation with the data you have : \[y = a \times x + b \rightarrow b = y + a \times x\] (with a : slope, and b : offset)

OpenStudy (anonymous):

lol.. thank you, but I'm still lost on how to get the slope. I suck at math royally. .. my work says to find the slope of the line and then find the linear equation in slope intercept form

OpenStudy (anonymous):

Ok, I'll give a numerical example with your data.

OpenStudy (anonymous):

Your data are : \[f(2) = 108000 \text { and } f(4) = 96000\] Plug the two in the slope equation : \[\frac{f(4) - f(2)}{4 - 2} \equiv \frac{96000 - 108000}{2} \equiv \frac{ -12000}{2} \equiv -6000 \] Your slope is -6000. The general equation of a linear function is : \[y = f(x) = a x + b \text{ and with your slope : } y = f(x) = -6000 x + b\] Solv ethe equation to find b, the offset at origin (don't know what you call it, that's how oit is named here in France) : \[y = -6000 x +b \equiv b = y + -6000x \text{ , so we fill with the data : }\] \[b = 96000 + -6000 \times 4 = 72000\] So the formula is : y = -6000x + 72000